# A Survey on the Applications of Zero-Knowledge Proofs

### v1

@@ III-B1 Motivation and Definition
Proof System(s) Intermediate Representation Syntax Comparison Purpose(s)
@@
Juvix Plonk, Halo2 VampIR Ocaml Anoma Interts
@@
TABLE II: Comparison of zkDSLs
III-B2 Methodology
Some zkDSLs that are listed are O1JS [38], Circom [23], Cairo [29], Noir [39], Juvix [40], and Lurk [41],.
The goal of a zkDSL is to translate a high-level language into an arithmetic circuit that can be passed into a proving system to output a proof on execution of a given program. Some of the most common circuits targeted are R1CS, Plonk, and AIR, as shown in the Arithmetization Schemes table 5. Each language has a niche that it is trying to fulfill. For example, Circom is a language that is focused on writing arithmetic circuits whereas Lurk is meant to be a generalized programming language.
While zkDSLs achieve the goal of abstraction, some drawbacks must be considered. First, zkDSLs encounter challenges in memory management complexity, primarily due to the inherent intricacies associated with zero-knowledge proofs and cryptographic computations. The nature of zkDSLs, which often involve complex mathematical transformations and intricate circuit representations, demands efficient memory management strategies. The need to handle cryptographic primitives, large-scale boolean circuits, and data structures introduces complexities that traditional memory management systems may find challenging to navigate. Languages like Cairo implement optimization techniques in their compiler or VM to reduce this overhead through constraint reduction and cycle elimination.
Figure 5: Arithmetization Schemes
Second, for some time, zkDSLs have a history of problems with the developer paradigms of recursion, conventional conditionals (excluding ternaries), mutable variables, and user-defined structures, causing unique coding challenges. Recursion, a staple in conventional programming, faces challenges due to the absence of a direct stack and the deterministic path required by circuit-based computation. The representation of if-else statements becomes intricate within algebraic circuits, favoring ternary expressions but sacrificing the straightforwardness of traditional conditionals. The immutability inherent in algebraic circuits complicates the emulation of mutable variables that change state during computation. Furthermore, the structured and hierarchical nature of user-defined structures, commonplace in languages like C++, presents a unique set of challenges. Developers navigating the ZKP circuit structure must consider these drawbacks, as the departure from traditional programming paradigms necessitates a reevaluation and adaptation of coding practices.
@@ III-B2 Methodology
III-B3 Applications
A classic example of a zkDSL is the circuit writing language Circom [23], which has been used in applications such as TornadoCash [42] or DarkForest [43]. Circom represents both a programming language and a compiler designed for the creation of arithmetic circuits that are subsequently compiled into Rank-1 Constraint Systems (R1CS). This intricate process allows programmers using Circom to articulate arithmetic circuits at a constraint level, with the compiler generating a file containing the R1CS description, as well as WebAssembly and C++ programs. These output programs facilitate the computation of all circuit values. Additionally, an open-source library, Circomlib [44], features circuit templates to be used in to support writing other circuits. On top of that, O1JS [38] is a library proficient in generating and validating Zero-Knowledge (ZK) proofs derived from R1CS. Collectively, this suite of software tools serves to abstract the intricacies of ZK proving mechanisms, providing an interface for modeling low-level descriptions of arithmetic circuits.
Lurk [41] is a Turing complete LISP-based programming language that autonomously generates zk-SNARKs for arbitrary programs. It presents programs as data to the universal Lurk interpreter circuit to achieve Turing completeness without compromising the size of the proof artifacts generated. This departure from traditional approaches aims to enhance the scalability and capabilities of a ZKP. Further, the need for ad-hoc compilation of programs into flat circuits, a conventional process that imposes significant constraints on the size and complexity of achievable computations, is eliminated.
Leo [45] is a programming language designed for the development of formally verified zero-knowledge applications on the blockchain. Leo establishes a execution environment devoid of restrictions on running time, stack size, or instruction sets, offering flexibility. Beyond the intrinsic benefits of ensuring application privacy and mitigating maximal-extractable value (MEV), Leo has two other helpful properties. Firstly, applications undergo formal verification concerning their high-level specifications. Secondly, the succinct verification of applications is made accessible to anyone, irrespective of the application’s size. Leo maintains a set of tools, including a testing framework, package registry, import resolver, remote compiler, formally defined language, and theorem prover, specifically tailored for general-purpose zero-knowledge applications.
Figure 7: Timeline of Popular zk-DSLs
To summarize, zkDSLs emerge as versatile tools with a broad spectrum of applications, ranging from smart contract programming to general computation. These specialized languages are instrumental in the development of privacy-preserving features within smart contracts, facilitating secure and confidential transactions on blockchain networks. Beyond blockchain, zkDSLs find utility in general computation scenarios, where their application extends to fields such as data privacy, machine learning, secure distributed computations, and confidential data analysis. In cryptographic applications and security protocols, zkDSLs play a pivotal role by enabling the creation of proofs for statements related to cryptographic properties, thereby enhancing the security and privacy of digital transactions and communication. This versatility positions zkDSLs as valuable tools for developers working on various applications, providing a user-friendly interface to harness the power of zero-knowledge proofs in diverse domains.
III-C Libraries and Frameworks
III-C1 Motivation and Definition
@@
In the domain of zero-knowledge proofs (ZKP), the development process often entails numerous operations characterized by a high barrier to entry, necessitating a profound understanding of computational strategies, cryptographic primitives, elliptic curves, and the intricacies of ZKP mechanisms. For developers, the hardship lies in crafting reusable code and mitigating redundancy, particularly within ZK proofs, where constraints such as circuit size and gas usage can impose significant limitations. To address these challenges, a multitude of libraries have emerged, offering implementations for cryptographic “gadgets” that facilitate the construction of Rank-1 Constraint Systems (R1CS) instances from modular “gadget” classes as listed in Table III. To faciliate the proving and verification process, elliptic curves as utilized, as mentioned in the previous section. Operations on these curves is relatively standardized and does not always need to be recreated, so libraries avoid the issue of duplicated code in this area. The combination of circuit gadgets and elliptic curves allow for the modular composition of an application in conjunction with a compatible proving system such as Groth16 [21] or Bulletproofs [22]. These libraries play a crucial role in abstracting the complexities of preprocessing in a modular manner, enabling developers to focus on higher-level design aspects and promoting code reusability across ZKP applications. This emphasis on modularization and abstraction enhances development efficiency and advances innovation within the ZKP domain by fostering collaboration and knowledge sharing among developers.
Circuit Gadgets
Arithmetic Operations
Booleans
Range Proofs
SHA256
MIMC
Poseidon
Pedersen
blake2
Lookups
Proving Systems
Groth16
Plonk
Halo
Bulletproofs
Pinocchio
Marlin
Gemini
Sonic
Brakedown
Spartan
Elliptic Curves
BLS12-381
BLS12-377
BN254
MNT-298
MNT-753
Jubjub
Secp256k1
Curve25519
BN256
TABLE III: Categories of Cryptographic Components
III-C2 Methodology
@@
In the design of ZKP frameworks and libraries, several fundamental principles are leveraged to optimize performance and efficiency. One crucial aspect is the operation over small fields, as demonstrated by systems like ethSTARK [46] and Plonky2 [47]. Unlike traditional elliptic curve groups, which require large field sizes (e.g., 256 bits) for standard security levels, these systems utilize smaller prime fields (e.g., 64 bits). This approach capitalizes on the efficiency of small-field arithmetic, resulting in state-of-the-art proving performance. Additionally, using small-field elements reduces storage requirements, improving CPUs’ cache efficiency. Another critical consideration is the flexibility in field selection. These schemes often use computationally structured fields, providing additional optimization opportunities. Finally, these frameworks and libraries prioritize using cheaper cryptographic primitives, ensuring cost-effective and efficient ZKP implementations [48].
Abstracting away finite field operations holds paramount importance in zero-knowledge proofs (ZKPs). This endeavor necessitates the implementation of fundamental arithmetic operations, including addition, subtraction, multiplication, (modular) exponentiation, and inverse exponentiation, over finite fields. Considering these operations are used so heavily in constructing a ZKP, their modularity and efficiency must be implemented to suffice for any use case. [49]
In elliptic curve cryptography, an elliptic curve is typically defined over a prime field of a specific order denoted as $F_{q}$. The elliptic curve group ($E(F_{q})$) comprises the subgroup of points within the field that lie on the curve, including a distinguished point at infinity. While some SNARKs function over elliptic curves without necessitating pairings, others rely on pairings and thus necessitate pairing-friendly elliptic curves. Pairings, a fundamental operation in cryptographic protocols, involve taking an element from the first group ($G_{1}$) and another element from the second group ($G_{2}$) to compute an element in the target group ($G_{T}$), typically denoted as $e(P,Q)$, where $P$ belongs to $G_{1}$ and $Q$ belongs to $G_{2}$. Efficient implementation of pairings, scalar multiplication, and multi-scalar multiplication (MSM) over pairing-friendly elliptic curves is essential for achieving computational efficiency. Therefore, prioritizing the efficient implementation of these operations is crucial for optimizing the performance of cryptographic protocols relying on pairing-friendly elliptic curves.[49]
To facilitate proof verification, ZKP frameworks often include the specification of cryptographic primitives within their design at the cost of the core SNARK implementation’s efficiency, as mentioned above. However, not all computations conform to arithmetic modulo $p$, necessitating the development of circuits for non-native field operations. For instance, verifying elliptic-curve-based cryptographic primitives, such as ECDSA signatures, requires computations over a different field, such as $\mathbb{Z}_{q}$, which SNARKs do not natively support. Additionally, SNARKs often exhibit inefficiencies when dealing with traditional hash algorithms and signature schemes, like SHA-2 and ECDSA. To address this, the community has proposed SNARK-friendly alternatives, such as Poseidon Hash, Keccak Hash, Pedersen Hash, MIMC Hash, and Ed25519 (EdDSA signature), which are specifically optimized for use within SNARKs. Despite the benefits of utilizing SNARK-optimized primitives, practical applications often necessitate CPU-optimized primitives due to constraints in non-native field arithmetic. For example, some applications require the verification of ECDSA signatures, which involves non-native field arithmetic, leading to numerous constraints that increase the complexity and verification time of the proof. [49]
By writing these building blocks piece by piece, researchers and developers can create SNARKs across diverse settings, thus facilitating informed decision-making and enhancing the practical utility of SNARK-based cryptographic protocols.
@@
TABLE IV: Popular Elliptic Curves in ZK Libraries
III-C3 Applications
Arkworks [50] is a Rust ecosystem tailored for zkSNARK programming, offering a set of libraries to streamline zkSNARK application development. These libraries implement essential components, including generic finite fields and R1CS constraints, enabling developers to integrate zkSNARK functionalities into their applications seamlessly. Arkworks also provides interfaces for various purposes, such as defining interfaces for SNARKs and relations (e.g., R1CS, AIR) and offering SNARK proving systems like Groth16 and Marlin. Further, the ecosystem includes tools for circuit building and algebraic operations for finite fields and elliptic curves, empowering developers to implement zkSNARKs in their projects efficiently.
Gnark [51] is a library that enables developers to design circuits using the programming language Go. It employs a versatile API and command line interface to accomplish the ZKP process in a familar way to developers. The two proving schemes that Gnark supports are Groth16 and Plonk, along with a variety of elliptic curves for developers to choose ranging such as BN254, BLS12-381, and BLS12-377. A core focus of the Gnark library is speed of the prover and verifier, since it offers an API for both the frontend and the backend of the proving process. The standard library of Gnark implements common functions such as the MiMC hash function, EdDSA signature verification, Merkle proof verification, and a generalized zk-SNARK verifier. The performance of circuits written in Gnark can be analyzed using the in suite profiling tools available within the library.
@@ III-C3 Applications
CirC [52] is a project dedicated to compiler infrastructure for cryptosystems and verification. It focuses on cryptographic tools such as proof systems, multi-party computation, and fully homomorphic encryption, typically applied to computations expressed as systems of arithmetic constraints. These applications require compilers that can translate high-level programming languages (e.g., C) into such constraints. CirC aims to provide a shared infrastructure for building constraint compilers, offering a valuable resource across various applications. These compilers can translate code into a Rank-1 Constraint System (R1CS), enabling efficient implementation of cryptographic protocols and verification mechanisms. Circify simplifies certain aspects of supporting a new language by transforming stateful programs with complex control flow into flat circuits. However, it does not address language-specific features like type-checking and namespacing. Circify supports various frontends, including C, ZoKrates [53] (Z#) and Circom. This flexibility allows developers to choose or build the frontend that best suits their needs while benefiting from Circify’s capabilities in managing program transformations for cryptographic applications.
The TypeScript library O1JS [38] is designed to cater to users with a web development background, offering a user-friendly approach to writing ZK programs and smart contracts for the Mina [54] blockchain. It is described as “an embedded DSL” and is executed as normal Typescript. This library permits developers to write arbitrary ZK programs utilizing many built-in provable operations, including basic arithmetic, hashing, signatures, boolean operations, and comparators. With o1js, developers can create zkApps on Mina, smart contracts that execute client-side and handle private inputs. The entire o1js framework is packaged as a single TypeScript library, making it accessible in web browsers and Node.js environments, thus allowing developers to integrate their ZK programs into existing web applications.
Various implementations of Zcash’s Halo2 [55] proving system, such as those by Axiom and the Ethereum Foundation, provide fundamental primitives for writing zero-knowledge proof circuits. The proving system involves several stages, from committing to polynomials that encode the main components of the circuit, including cell assignments, permuted values, products for lookup arguments, and equality constraint permutations. The next step is constructing the vanishing argument, which constrains all circuit relations to zero, including standard and custom gates, lookup argument rules, and equality constraint permutation rules. The polynomials are then evaluated at all necessary points, including relative rotations used by custom gates and vanishing argument pieces. Finally, the multipoint opening argument is constructed to ensure the consistency of evaluations with their commitments, and the inner product argument is run to create a polynomial commitment opening proof for the multipoint opening argument polynomial. Halo2 is known as a relatively low-level library with high customizability.
Nova [56] is a recursive SNARK that enables incrementally verifiable computation (IVC), a cryptographic primitive that allows a prover to produce a proof of correct execution of a “long-running” sequential computation in an incremental fashion. IVC allows for proofs to build on top of each other in an efficient fashion that speeds up the entire proving and verification process. Nova is implemented in Rust and supports three curve cycles: Pallas/Vesta, BN254/Grumpkin, and secp/secq. It supports frontends such as Bellpepper [57], and its native APIs accept circuits expressed with Bellpepper, Circom, and Lurk.
III-D Hardware Acceleration
III-D1 Motivation and Definition
@@ III-D2 Methodology
Hardware acceleration is defined as the use of optimizing or creating dedicated computer components to improve the performance and efficiency of a specific operation. The main instruments used for this acceleration such as field programmable gate arrays (FPGAs), graphics processing units (GPUs), and application-specific integrated circuits (ASICs) in the ZK world. The limiting factors of hardware acceleration projects are the memory capacity, speed of memory access, speed of data transfer, and speed of arithmetic units. In proof systems where both number theoretic transforms (NTTs) and multi-scalar multiplications (MSMs) are used, the majority of the proof generation time is spent on MSMs, with NTTs accounting for the remainder. Both MSMs and NTTs present performance challenges that can be addressed in several ways. MSMs can be executed on multiple threads, allowing for parallel processing. However, when dealing with large vectors of elements, the multiplications may still be slow and demand considerable memory resources. Additionally, MSMs face scalability issues and can remain sluggish even when extensively parallelized. On the other hand, NTTs involve frequent data shuffling during the algorithm, making them difficult to distribute across a computing cluster. They also require significant bandwidth when run on hardware due to the need to load and unload elements from large datasets. For instance, if a hardware chip has 16GB of memory or less, running NTTs on a dataset larger than 100GB would necessitate data transfers over the network, significantly slowing down the operations [58].

### v2

@@ 3.2 Domain Specific Languages
zkDSL Proof System(s) Intermediate Representation Syntax Comparison Purpose(s)
@@
Juvix Plonk, Halo2 VampIR Ocaml Anoma Intents
@@
3.2.1 Motivation and Definition
zkDSL (zero-knowledge domain-specific language) bridges high-level program representation and the low-level details of ZKPs, like writing circuits. Its purpose is to convert high-level language into an arithmetic circuit, which can then be used in a proving system to produce a proof for a program’s execution. Common circuit targets include R1CS, Plonk, and AIR. Each zkDSL has a specific focus; for instance, Circom [20] specializes in writing arithmetic circuits, while Lurk [6] is a general-purpose programming language. Table 4 gives a mapping from language frontends to the constraint targets and proof backends they commonly support.
3.2.2 Methodology
While zkDSLs offer abstraction, they have some drawbacks. They face memory management challenges due to the complexities of zero-knowledge proofs and cryptographic computations, requiring efficient handling of cryptographic primitives, large boolean circuits, and data structures. Additionally, zkDSLs struggle with conventional programming paradigms, such as recursion, traditional conditionals, mutable variables, and user-defined structures. In practice, circuit and DSL ecosystems must defend against several common classes of bugs: (i) under-constrained circuits, where the constraint system does not fully enforce the intended relation and allows invalid witnesses; (ii) unsafe gadgets, where reusable components omit edge cases for example, missing constraints in arithmetic or comparator gadgets; (iii) range-check mistakes, where field elements are not properly bounded and can represent unintended values; and (iv) witness malleability patterns, where multiple witnesses satisfy the same public statement in unintended ways. To mitigate these risks, projects increasingly use a mix of unit tests, property-based testing, fuzzing, differential testing between backends, and audit-driven checklists. Practitioner resources include public audit reports and issue trackers for major proving systems and toolchains, which provide concrete examples of real-world circuit failures and patch patterns. Figure 4 provides a compact view of how zkDSLs translate developer code into constraints and then into proofs. The main takeaway is that performance and security depend on both compilation choices (constraint generation) and backend choices (proving and commitment schemes).
CodeConstraint SystemCryptographicCompilerProving SystemProof
@@ 3.2.2 Methodology
3.2.3 Projects
To organize the zkDSL ecosystem more clearly, we group representative projects into three categories: (i) constraint-centric DSLs that prioritize explicit circuit construction (Circom), (ii) higher-level languages and VM-oriented stacks that compile programs into an execution trace or constraint IR (Cairo and Lurk), and (iii) compiler frameworks that target multiple constraint backends (CirC). We use these categories when relating tools to applications such as rollups, privacy protocols, and ZKML. Circom [20] has been used in applications such as TornadoCash [75] and DarkForest [1]. Circom is both a programming language and a compiler for the creation of arithmetic circuits, which are subsequently compiled into Rank-1 Constraint Systems (R1CS). Circomlib [44], an open-source library, provides reusable circuit templates. O1JS [61] generates and validates zero-knowledge proofs from R1CS. Together, these tools abstract the low-level details of ZK proving, providing a higher-level interface for circuit construction.
Lurk [6] is a Turing complete LISP-based programming language that autonomously generates zkSNARKS for arbitrary programs. It presents programs as data to the universal Lurk interpreter circuit to achieve Turing completeness without compromising the size of the proof artifacts generated. Leo [19] is a programming language for developing formally verified zero-knowledge applications on the blockchain. It offers an unrestricted execution environment with no limits on running time, stack size, or instruction sets, providing flexibility. Besides ensuring application privacy and reducing maximal-extractable value (MEV), Leo has two key features: formal verification of applications against high-level specifications and accessible succinct verification regardless of application size.
3.3 Libraries and Frameworks
3.3.1 Motivation and Definition
@@
In the domain of zero-knowledge proofs (ZKP), the development process has a high barrier to entry, requiring working knowledge of computational strategies, cryptographic primitives, elliptic curves, and ZKP protocol design. For developers, the hardship lies in crafting reusable code and mitigating redundancy, particularly within ZK proofs, where constraints such as circuit size and gas usage can impose significant limitations. To address these challenges, a multitude of libraries have emerged, offering implementations for cryptographic gadgets that facilitate the construction of Rank-1 Constraint Systems (R1CS) instances from modular gadget classes as listed in Figure 5. To facilitate the proving and verification process, elliptic curves are utilized. Operations on these curves is relatively standardized and does not always need to be recreated, so libraries avoid the issue of duplicated code in this area. The combination of circuit gadgets and elliptic curves allow for the modular composition of an application in conjunction with a compatible proving system such as Groth16 [41] or Bulletproofs [17].
These libraries abstract the complexities of preprocessing in a modular manner, letting developers focus on higher-level design and reuse code across ZKP applications. A central concept used throughout libraries and backends is a commitment. A commitment is a cryptographic binding to a value that can be opened later. Informally, it provides binding (the committer cannot change the committed value after committing) and often hiding (the commitment does not reveal the value). In ZK systems, commitments are used to bind the prover to large objects that the verifier cannot read in full, such as polynomials, Merkle roots, or witness encodings. This is why “Commitments” appears as an explicit input in Figure 3: the prover commits to artifacts that will later be checked via small openings rather than full disclosure.
Figure 5: Categories of Cryptographic Components
3.3.2 Methodology
@@
In the design of ZKP frameworks and libraries, several fundamental principles are leveraged to optimize performance and efficiency. One important design choice is the operation over small fields, as demonstrated by systems like ethSTARK [94] and Plonky2 [79]. Unlike traditional elliptic curve groups, which require large field sizes (e.g., 256 bits) for standard security levels, these systems utilize smaller prime fields (e.g., 64 bits). This approach capitalizes on the efficiency of small-field arithmetic, resulting in state-of-the-art proving performance. Additionally, using small-field elements reduces storage requirements, improving CPUs’ cache efficiency. Another critical consideration is the flexibility in field selection. These schemes often use computationally structured fields, providing additional optimization opportunities. Finally, these frameworks and libraries prioritize using cheaper cryptographic primitives, ensuring cost-effective and efficient ZKP implementations [64].
Abstracting away finite field operations necessitates the implementation of fundamental arithmetic operations, including addition, subtraction, multiplication, (modular) exponentiation, and inverse exponentiation, over finite fields in a modular and efficient manner. Efficient implementation of pairings, scalar multiplication, and multi-scalar multiplication (MSM) over pairing-friendly elliptic curves is also essential for achieving computational efficiency. Efficient implementation of these operations directly affects the performance of cryptographic protocols that rely on pairing-friendly elliptic curves. SNARKs can be inefficient with traditional hash algorithms and signature schemes, like SHA-2 and ECDSA, for example due to non-native field arithmetic. To address this, SNARK-friendly alternatives such as Poseidon Hash, Keccak Hash, Pedersen Hash, MIMC Hash, and Ed25519 (EdDSA signature) have been proposed, optimized for SNARKs [26].
Concrete optimization themes that repeatedly yield large speedups in practice include: (i) MSM improvements (for example, Pippenger-style bucket methods, endomorphism-based speedups, and better batching), (ii) FFT/NTT tuning (for example, domain selection, parallelization, and cache-aware layouts), (iii) lookup arguments and custom gates that reduce constraint counts for common primitives, (iv) recursion and proof composition that amortize verification across many steps (for example, rollups and IVC), and (v) hardware acceleration that targets MSM/NTT/hashing kernels.
Table 5: Popular Elliptic Curves in ZK Libraries
@@
Table 5 focuses on elliptic-curve support that is central to many SNARK libraries. In contrast, STARK-style systems often rely on hash-based polynomial commitments and Merkle-based authentication rather than pairing-friendly curves, which is one reason their setup models and performance profiles differ from pairing-based SNARK families. A recursive SNARK is a proof system where the statement being proved includes the correctness of verifying another proof. Concretely, recursion proves that a verifier circuit accepts a previous proof together with a new computation step, allowing many steps to be composed into a single succinct proof. This enables incremental verifiable computation (IVC) and constant-size proofs for long-running computations, and it is widely used in rollups, proof-carrying data, and systems that require repeated updates with bounded verification cost.
3.3.3 Projects
@@
Arkworks [7] is a Rust ecosystem tailored for zkSNARK programming, offering a set of libraries to streamline zkSNARK application development. These libraries implement essential components, including generic finite fields and R1CS constraints, enabling developers to seamlessly integrate zkSNARK functionality into their applications. Gnark [15] is a library that enables developers to design circuits in Go. It employs a versatile API and command line interface to accomplish the ZKP process in a familiar way to developers. The two proving schemes that Gnark supports are Groth16 and Plonk, along with a variety of elliptic curves for developers to choose from, such as BN254, BLS12-381, and BLS12-377. A core focus of the Gnark library is speed of the prover and verifier, since it offers an API for both the frontend and the backend of the proving process. CirC [73] is a project dedicated to compiler infrastructure for cryptosystems and verification. It focuses on cryptographic tools such as proof systems, multi-party computation, and fully homomorphic encryption, typically applied to computations expressed as systems of arithmetic constraints. These applications require compilers that can translate high-level programming languages (e.g., C) into such constraints. CirC aims to provide a shared infrastructure for building constraint compilers, offering a valuable resource across various applications. These compilers can translate code into a Rank-1 Constraint System (R1CS), enabling efficient implementation of cryptographic protocols and verification mechanisms.
The TypeScript library O1JS [61] is designed to cater to users with a web development background, offering a user-friendly approach to writing ZK programs and smart contracts for the Mina [60] blockchain. It is described as an embedded DSL and is executed as normal TypeScript. This library permits developers to write arbitrary ZK programs utilizing many built-in provable operations, including basic arithmetic, hashing, signatures, boolean operations, and comparators. Various implementations of Zcash’s Halo2 [25] proving system, such as those by Axiom and the Ethereum Foundation, provide fundamental primitives for writing zero-knowledge proof circuits. The proving system involves several stages, from committing to polynomials that encode the main components of the circuit, including cell assignments, permuted values, products for lookup arguments, and equality constraint permutations. Nova [51] is a recursive SNARK that enables incrementally verifiable computation (IVC), a cryptographic primitive that allows a prover to produce a proof of the correct execution of a long-running sequential computation incrementally. IVC allows proofs to build on each other efficiently, speeding up the entire proving and verification process.
3.4 Hardware Acceleration
3.4.1 Motivation and Definition
@@ 3.4.2 Methodology
Hardware acceleration is the use of optimized or dedicated computer components to improve the performance and efficiency of a specific operation. The main instruments used to accelerate this in the ZK world are field-programmable gate arrays (FPGAs), graphics processing units (GPUs), and application-specific integrated circuits (ASICs). The limiting factors for hardware acceleration projects are memory capacity, memory access speed, data transfer speed, and arithmetic unit speed. In proof systems where both number theoretic transforms (NTTs) and multi-scalar multiplications (MSMs) are used, the majority of the proof generation time is spent on MSMs, with NTTs accounting for the remainder. Both MSMs and NTTs present performance challenges that can be addressed in several ways. MSMs can be executed across multiple threads, enabling parallel processing. However, when performing large vector operations, multiplications may still be slow and require considerable memory. Additionally, MSMs face scalability issues and can remain sluggish even when extensively parallelized. On the other hand, NTTs involve frequent data shuffling during the algorithm, making them difficult to distribute across a computing cluster. They also require significant bandwidth when run on hardware due to the need to load and unload elements from large datasets.